# A hidden Markov model for statistical arbitrage in international crude oil futures markets

### v2

@@ 4.4. Robustness analysis
To assess the robustness of our findings, we now systematically repeat the analysis carried out in the previous sections changing the transaction costs parameter $c$, the starting date $t_{0}$ of the sample and the breaking date $t_{B}$ that divides the training sample from the test one.
@@ 4.4.1. Sensitivity with respect to transaction costs
Figure 9. Performances of the strategies with respect to $c$; $t_{0}=$ 03/26/2018, $t_{B}=$ 07/01/2022.
Figure 9 displays the performances of the statistical arbitrage strategies as a function of $c$. Obviously, both the annualized return $R$ and the Sharpe ratio $SR$ are declining in $c$. The slope of this decline is proportional to the frequency of opening/closing positions: the performances of strategies PV, PredI and PI that prescribe the opening of several trades (see Figure 6) are the most sensible to $c$ while strategies ProbI and RI are less affected by transaction costs. We find that, over the considered time period, statistical arbitrage strategies that include the Shanghai futures are quite robust with respect to the level of transaction costs. In particular, the benchmark S&P passive strategy becomes preferable only when $c$ approaches 1%. It is also noteworthy that the strategies that make use of the OU-HMM model introduced in Section 2.1, namely the PredI and the PI strategies, turn out to be very profitable for small (but realistic) levels of transaction costs. In our view, this represents a further evidence of the potential advantages of relying on a stochastic model in the design of statistical arbitrage strategies.
@@ 4.4.2. Sensitivity with respect to $t_{0}$
Figure 10. Performances of the strategies with respect to $t_{0}$; $c=80$bps, $t_{B}=$ 07/01/2022. Missing datapoints are due to the rejection of cointegration among the futures contract over the related time window.
Figure 10 depicts the performance of our strategies as a function of the starting date $t_{0}$ of the sample. Since the breaking date $t_{B}$ is fixed, if $t_{0}$ moves forward the number of observations in the training sample shrinks. This is expected to worsen the quality of the filter-based estimates derived according to the algorithm described in Section 2.2, due to the availability of a smaller number of observations. As we can see from Figure 10, reducing the training sample weakens the cointegrating relationship among the time series and, as a consequence, lowers the performance of our strategies. This indicates that, in order to be profitable, our statistical arbitrage strategies require a minimum sample size for the model to be accurately estimated. From Figure 10 we can also notice that the dataset is significantly affected by the market turmoil of the COVID-19 pandemics. Given the rather recent introduction of the Shanghai futures, this issue cannot be avoided, since the available dataset is relatively small.
@@ 4.4.3. Sensitivity with respect to $t_{B}$
Figure 11. Performances of the strategies with respect to $t_{B}$; $c=80$bps, $t_{0}=$ 03/26/2018.
Figure 11 shows the performance of our strategies as a function of the breaking date $t_{B}$. As $t_{B}$ moves backward, the training sample shrinks, while the test sample becomes longer. The impact on the performance of statistical arbitrage strategies is therefore twofold. On the one hand, reducing the size of the training sample weakens the reliability of estimated cointegration model. On the other hand, a larger test sample provides more opportunities for profitable trades. In our dataset, a small decline in the performance is registered for $t_{B}$ before than March-July 2022. Bringing $t_{B}$ forward has ambiguous effects as well: the cointegration is estimated over a larger training sample but there is less time to take advantage of that. As shown in Figure 11, in our dataset we document a small increase of the performances for $t_{B}$ after March-July 2022.
@@ 5. Conclusions
In this work, we have investigated statistical arbitrage strategies that generalize pairs trading strategies, involving the two established crude oil futures contracts, the Brent and the WTI, together with the more recently introduced Shanghai crude oil futures. We have documented that the time series of the three futures prices are cointegrated. In order to dynamically model the resulting cointegration spread, we have proposed a mean-reverting stochastic process with regime switching modulated by an unobservable Markov chain. The model is estimated by relying on a filter-based version of the EM algorithm, thereby ensuring that the model stays dynamically tuned to the market situation. We have considered and analyzed several statistical arbitrage strategies, showing that strategies designed by relying on our stochastic model are remarkably profitable, even under conservative levels of transaction costs. Based on these results, we believe that it would be interesting to evaluate alternative investment strategies based on our stochastic model and study optimal portfolio problems by relying on the techniques of stochastic optimal control under partial information.
Our findings suggest that, despite the well established cointegration among the Brent, the WTI and the Dubai crude oil futures, profitable statistical arbitrages in these three assets seem to have been exhausted. This is not the case for the Shanghai futures, which yields significant investment performances even if subject to a non-negligible level of risk. The introduction of the Shanghai futures is quite recent and, therefore, more extensive empirical analysis will be needed in order to assess the specific features of this security in the context of global financial markets.

### v3

@@ 4.4. Robustness analysis
To further assess the robustness of our findings, we now systematically repeat the analysis carried out in the previous sections for different values of the transaction costs parameters $c_{i}$’s, the starting date $t_{0}$ of the sample and the breaking date $t_{B}$ that separates the training sample from the test one. Similarly as in Section 4.2.1, the analysis is performed using the baseline bandwidth level $\alpha=0.20$, representing an balanced specification among the range of $\alpha$ considered in Section 4.2.
@@ 4.4.1. Sensitivity with respect to transaction costs
Figure 8. Performances of the strategies with respect to different percentages of the vector collecting the original transaction costs $c_{i}$ in Table 1; $t_{0}=$ 03/26/2018, $t_{B}=$ 07/01/2022.
Figure 8 displays the performances of the statistical arbitrage strategies as a function of a multiple of the original values of the transaction cost parameters in Table 1. Obviously, both the annualized return $R$ and the Sharpe ratio $SR$ are declining in $c$. The slope of this decline is proportional to the frequency of opening/closing positions: the performances of strategies PV, PredI and PI that prescribe the opening of several trades (see Figure 6) are the most sensible to the $c$’s while strategies ProbI and RI are less affected by transaction costs. We find that, over the considered time period, statistical arbitrage strategies that include the Shanghai futures are quite robust with respect to the level of transaction costs. It is also noteworthy that the strategies that make use of the AR-HMM model introduced in Section 2.1, namely the PredI and the PI strategies, turn out to be very profitable for small (but realistic) levels of transaction costs. In our view, this represents a further evidence of the potential advantages of relying on a stochastic model in the design of statistical arbitrage strategies.
@@ 4.4.2. Sensitivity with respect to $t_{0}$
Figure 9. Performances of the strategies with respect to $t_{0}$; $t_{B}=$ 07/01/2022. Missing datapoints are due to the rejection of cointegration among the futures contract over the related time window.
Figure 9 depicts the performance of our strategies as a function of the starting date $t_{0}$ of the sample. Since the breaking date $t_{B}$ is fixed, if $t_{0}$ moves forward the number of observations in the training sample shrinks. This is expected to worsen the quality of the filter-based estimates derived according to the algorithm described in Section 2.2, due to the availability of a smaller number of observations. As we can see from Figure 9, reducing the training sample weakens the cointegrating relationship among the time series and, as a consequence, lowers the overall performance of our strategies. This indicates that, in order to be profitable, our statistical arbitrage strategies require a minimum sample size for the model to be accurately estimated. From Figure 9 we can also notice that the dataset is significantly affected by the market turmoil of the COVID-19 pandemics. Given the rather recent introduction of the Shanghai futures, this issue cannot be avoided, since the available dataset is relatively small.
@@ 4.4.3. Sensitivity with respect to $t_{B}$
Figure 10. Performances of the strategies with respect to $t_{B}$; $t_{0}=$ 03/26/2018.
Figure 10 shows the performance of our strategies as a function of the breaking date $t_{B}$. As $t_{B}$ moves backward, the training sample shrinks, while the test sample becomes longer. The impact on the performance of statistical arbitrage strategies is therefore twofold. On the one hand, reducing the size of the training sample weakens the reliability of estimated cointegration model. On the other hand, a larger test sample provides more opportunities for profitable trades. In our dataset, a small decline in the performance is registered for $t_{B}$ after January 2022. Bringing $t_{B}$ forward has ambiguous effects as well: the cointegration is estimated over a larger training sample but there is less time to take advantage of that. As shown in Figure 10, in our dataset we document a small increase of the performances for $t_{B}$ after July 2022.
@@ 5. Conclusions
In this work, we have investigated statistical arbitrage strategies that extend classical pairs trading strategies, involving the two established crude oil futures contracts, the Brent and the WTI, together with the more recently introduced Shanghai crude oil futures. We have documented that the time series of the three futures prices are cointegrated. In order to dynamically model the resulting cointegration spread, we have proposed a mean-reverting process with regime switching modulated by an unobservable Markov chain. The model is estimated by relying on a filter-based version of the EM algorithm, thereby ensuring that the model stays dynamically tuned to the market situation. We evaluate several statistical arbitrage strategies based on both backward-looking and model-based forward-looking signals. The empirical results show that strategies grounded in the proposed regime-switching model deliver economically meaningful and robust performance, even under conservative levels of transaction costs. In particular, incorporating the Shanghai crude oil futures enhances profitability, suggesting that this relatively new asset yields relative-value trading opportunities. Moreover, the use of more than two futures contracts enables effective arbitrage execution even when pairwise cointegration is absent.
Based on these results, we believe that it would be interesting to evaluate alternative investment strategies based on our stochastic model and study optimal portfolio problems by relying on the techniques of stochastic optimal control under partial information.